The reaction rate of CO and NO2 in the reaction

CO(g) + NO2(g) → CO2(g) + NO(g)

is measured using the initial rates method. The results are tabulated below.

[CO] (mol/L)

NO2 (mol/L)

-([CO]/Δt (mol/L·s)

8.00 10-4

5.50 10-4

8.40 10-8

8.00 10-4

7.78 10-4

1.68 10-7

1.60 10-3

5.50 10-4

1.68 10-7

Determine the rate expression and calculate the rate constant for the reaction.
Answer #1

Solution:

Part A) For the given reaction,

CO + NO2 = CO2 + NO

In the given data,

From trial 1 and 2, when [CO] is constant, then on increasing nealy 50 % the concentration of NO2, the rate is increased upto 100 %

Hence, order with respect to NO2 = 1/2

From trial 2 and 3, when concentration of NO2 is constant, then on douling the concentration of CO increases rate two times.

Thus, order with respect to CO = 1

Therefore, rate expression will be:

Rate = k [NO2]^1/2  [CO]

Part B) Calculation of rate constant:

From first trial, putting the values in above rate law,

Rate = k [NO2]^1/2  [CO]

8.4 x 10^-8  M s-1 = k x (5.50 x 10^-4 M)^1/2 (8.0 x 10^-4 M)

k = 8.4 x 10^-8 M s-1 / 2.35 x 10^-2 M^(-1/2) x 8.0 x 10^-4 M

k = 4.46 x 10 ^-3 M^1/2 s-1